A 50 kg skydiver loses 2400 kJ of energy during a jump. What was her change in height?


1
Expert's answer
2021-02-06T16:42:09-0500
"PE_{loss}=mg\Delta h,""\Delta h=\dfrac{PE_{loss}}{mg}=\dfrac{2.4\cdot10^6\ J}{50\ kg\cdot9.8\ \dfrac{m}{s^2}}=4900\ m."